8-pack control system working group meeting

4 March 2003

In attendance: Chestnut, Hall, Kim, Nelson, S Smith

Notes by: J Nelson



-Database work is done, the triggers look ok.

-Need to check out readback with real signals, check archiving of variables (pwr, pmin, pmax, etc), and check the plots

-The confusion with LD1 and LD2 has been resolved.



-All admired and approved Judy’s panel with the attenuation and pulse width control



DAC/ADC Issues

All of these issues are for later:


-Janice gave panels to look at to Ron for which phase PVs we care about.  He may work on this sooner.  TBD.  Previously it was decided that one could input any number but then it’d be given the equivalent value between +/- 180.

-TBD: The trickle down of the SLED variables.  This was specified at a previous meeting (numbers on SLED panel need to trickle down to I/Q and A/P PVs).  For processing, we might want this feature functional sooner rather than later. 

-TBD: the trickle down of the variables on the power distribution panel as well.  There are 4 PVs there called Klystron Sled-in phases.  These need to trickle down to the phase DACs as well.  This phase is to be added to the channel phase angle (IQ calib panel) and to the timestep phase before going out to the DAC.



-This is on hold for now:  SWE will make a PV for a temporary DIO channel as a reset to Joe Frisch’s kludge RE Interlock.  They will send the name to Janice who will put a button on the panel with the pulse width and the klystron attenuation.  Its DIO channel assignment should be on Andrew’s previously sent list.  Also need a channel for status when it trips off.

-(no change) The following item was deferred for a few weeks:

LLRF + NLCTA Ops need to sort out what the LLRF enable and the mod off buttons do


SCP Configs

-Janice reported all TRIGs in crates 2 & 3 have been changed to TRBRs.  Steve reported that they’re working fine.

-Janice will make mini-dbedit files for Steve and Andrew to use to change the TRBRs’ rep rate. 



-Software ready to go.  Need to check out with real signals: PVs, plots, and archiving

-there are now 8 installed in the crate (6 from before, 2 recently).  Kukhee reported the 6th of the 6 has the wrong trigger rate and the 2 recently installed need clock/trig setup, which Steve will provide.

-(no change) We deferred the following: longer term issue: what to do if reading out an SIS when a fault occurs?

-(no change) Right now the SISs are paused and read out with a user request only.  A read when a fault happens will also be added to this.



-Janice reported the selection criteria looked good. 

-Bob will implement Judy’s most recently generated list of PVs to archive.

-Janice will check the archiving of the waveforms and perhaps with Steve/Andrew’s help send in a fake signal.



-Bob is working on Matlab functions to allow users to access the archived waveforms.  See previous notes for specs.  Bob reported good progress with more work to do.

-(no change) SWE plans to implement the following two features requested by LLRF (Joe’s email):

have waveforms archived on (1) every trip and (2) when an acquire is done by a user

-(no change) Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)



-All OKOK. 



(No change) All of the following will get going in the next few weeks:

-Each IQA channel has 6 or 7 calibration numbers that are module dependent.  Gain/offset I,Q, quadrature angle and overall phase.  These numbers are used to change the SIS waveforms from I/Q to A/P.  These numbers will go on the IQA module panel. 

-Steve will email the equations for this calculation

-Kukhee needs to change the SIS converter subrecord and new PVs need to be created.

-Once names are decided upon, Janice can make the panel (Judy may have started this already.)

-These 6 (or 7) numbers x 8 channels x 4 IQAs need to be archived at least once every 10 minutes.  The panel interface could just be one table for each IQA with 8 rows for the 8 channels and 6 columns with the calibration variables.


Processing algorithm

- Ron and Kukhee reported that the work is ongoing, the processing loop should be ready for testing in the next 2-3 weeks. 

- 10 powers were chosen for the processing loop to check – a reflected power for each klystron (4), a reflected power for each load (4), and a missing energy for each pair of klystrons (2) defined as the sum of the output forward power of the two klystrons in the pair minus the sum of the forward power to the two loads.  The PV names for these are:




And for the missing energy calculation:




Steph named the lost power PVs: TRS8:RF(6,8):LOST:PWR

-The processing loop is being written to change the pulse width in ns and the attenuation in dB.  Steve will provide a conversion from DAC volts to these units in the near future.

-Using the vacuum status (stat.sevr or something similar) is fine.  Right now that summary also checks the PLC status (which is currently red).  If this becomes a problem, we will either bypass that interlock, or ask Judy to take the PLC status out of the summary status.


Peak Power Meter (PPM)

-On hold until Tonee’s return.

-The descriptive string is being written to errlog when changed.

-(no change) Tonee will tweak the labview program to send the waveform to epics once we have names – probably late next week

-(no change) Tonee and Janice will set up the necessary hardware in ESB (Enet box, cables, etc)

-Judy made the panel which looks great.



-(no change) We’re currently waiting for Dick’s setup to settle down then we’ll do one more iteration of getting list, changing database, changing panels, programming the dcm, changing alarm handler and archiver stuff as needed



-Steve will try to use Ron’s workaround for getting epics to printout on the color printer near his office.  Either with the vax queue name (SLC:Qname) or the NT queue name (something similar).

-Judy is still waiting for the official queue name so the workaround isn’t needed.

-We looked at the A&P plots for klystron 5 to show the lovely changes Judy made so the klystron number appears at the top of each plot.  However, we noticed the phase plots have vertical limits of +/- 100 degrees instead of +/- 180.  Janice will fix.

-(no change)Judy reported the crate profiles on the web are updated and look good.  Janice requested that channel assignment information also be available via the web, even it it’s just a link to the template file.  Judy and Ron will look into this.

-Andrew will tell Kukhee when he takes out the DIO module to modify it for the deadman.  (no change: Janice requested that a bypass be incorporated into the design)

-(no change) we deferred the following:

-public/private matlab issues to be resolved at a later date

-LLRF needs to decide what the reset button does.



Sort task list by person


Ron (and otherwise unassigned SWE tasks):

-(on hold for now)PV/database for DIO reset of Joe Frisch’s kludge RE Interlock

-phase wraparound

-trickle down of sled variables

-trickle down of phase variables

-database/PVs for 6 channels of calibration numbers for each IQA channel (also connect to Kukhee’s SIS converter subroutine)

-processing algorithm



-Processing algorithm



-mini-dbedit files for changing TRBR rates

-checkout of waveform archiving, ADC readbacks, etc

-panel for 6 calib numbers for each IQA channel

-PPM panel, work with Tonee on PPM labview and hardware

-fix the A&P dm2k plot so phase is +/- 180 degrees.



-check into making module channel assignments available on the web as well as the crate profiles



-pass on equation for the 6 calibration numbers for each IQA channel

-kludge interlock reset hardware



-matlab functions to fetch archived waveforms.